# HIP 65426 is a High-Frequency Delta Scuti Pulsator in Plausible Spin-Orbit Alignment with its Directly Imaged Exoplanet

### v1

@@ Abstract
HIP 65426 hosts a young giant planet that has become the first exoplanet directly imaged with JWST. Using time-series photometry from the Transiting Exoplanet Survey Satellite (TESS), we classify HIP 65426 as a high-frequency Scuti pulsator with a possible large frequency separation of 7.230.02 cycles day. We check the TESS data for pulsation timing variations and use the nondetection to estimate a 95% dynamical mass upper limit of 12.8 for HIP 65426 b. We also identify a low-frequency region of signal that we interpret as stellar latitudinal differential rotation with two rapid periods of 7.850.08 hr and 6.670.04 hr. We use our TESS rotation periods together with published values of radius and to jointly measure the inclination of HIP 65426 to . Our stellar inclination is consistent with the orbital inclination of HIP 65426 b () at the percent level based on our orbit fit using published relative astrometry. The lack of significant evidence for spin-orbit misalignment in the HIP 65426 system supports an emerging trend consistent with preferential alignment between imaged long-period giant planets and their host stars.
@@ 1 Introduction
HIP 65426 (HD 116434, TIC 438702139) is a nearby (107.50.4 pc, Gaia Collaboration et al. 2016, 2021) A2 V star (Houk, 1978) that hosts a directly imaged giant planet (HIP 65426 b, Chauvin et al., 2017). The host star is a member of the Lower Centaurus-Crux (LCC) young moving group (de Zeeuw et al., 1999; Luhman, 2022), with a nominal age of 10–23 Myr (e.g., Mamajek et al., 2002; Sartori et al., 2003; Pecaut et al., 2012; Song et al., 2012). Independent of its membership in LCC, Chauvin et al. (2017) derived an age of 144 Myr for HIP 65426 primarily using isochrone fitting. HIP 65426 has several astrophysically interesting qualities in addition to being an exoplanet host star. Its (261–299 km s, Chauvin et al. 2017; Petrus et al. 2021) indicates that it is rapidly rotating. Hints of stellar pulsations have been identified in its radial velocity variations (e.g., Chauvin et al., 2017; Grandjean et al., 2020; Petrus et al., 2021), but HIP 65426’s variability status has not yet been formally confirmed and classified. A detection of high-frequency Scuti pulsations would independently confirm HIP 65426’s youth (Bedding et al., 2020) and might even enable an asteroseismic age to be measured (e.g., Murphy et al., 2021; Steindl et al., 2022).
The giant planet HIP 65426 b orbits at a semi-major axis of 62–120 au and inclined 99–112 to the line of sight based on orbit fitting analyses (Cheetham et al., 2019; Bowler et al., 2020; Carter et al., 2023; Do Ó et al., 2023; Blunt et al., 2023). Mass estimates of HIP 65426 b, derived from evolution models and from spectral fitting, span 6–11 (Chauvin et al., 2017; Cheetham et al., 2019; Marleau et al., 2019; Petrus et al., 2021; Carter et al., 2023). HIP 65426 was recently observed with NIRCam and MIRI as part of the JWST (Gardner et al., 2006, 2023) Early Release Science Program (Hinkley et al., 2022), resulting in HIP 65426 b becoming the first direct detection of an exoplanet at wavelengths beyond 5 m (Carter et al., 2023).
@@ 2 TESS Observations
HIP 65426 was observed by TESS in 2-minute cadence mode for Sectors 11, 38, and 64. The gaps between these sectors are 2 yr. Sector 11 observations spanned UTC 23 April 2019–20 May 2019, Sector 38 spanned UTC 29 April 2021–26 May 2021, and Sector 64 spanned UTC 06 April 2023–04 May 2023. We used lightkurve (Lightkurve Collaboration et al., 2018) to download PDC-SAP light curves processed by SPOC (Smith et al., 2012; Stumpe et al., 2012, 2014; Jenkins et al., 2016). Removing outlying photometry via a sigma-clipping of five standard deviations from the median flux value resulted in 16,392 cadences for Sector 11, 18,486 cadences for Sector 38, and 18,849 cadences for Sector 64.
We assessed our TESS data for any significant sources of contamination. We began by inspecting a region within a radius of 80″ from HIP 65426, which is an approximate size of the default aperture used for our light curves. We identified no Gaia sources with comparable brightness to HIP 65426, which has an apparent Gaia -band magnitude of 7.00 (Gaia Collaboration et al., 2016, 2023). The five brightest Gaia sources after HIP 65426 in this region (spanning seperations of 17–47″) have respective apparent -band magnitudes of 11.98, 13.28, 14.12, 14.59, and 15.05. HIP 65426 has no known stellar companions (Kouwenhoven et al., 2005; Chauvin et al., 2017; Petrus et al., 2021). Furthermore, the TESS Input Catalog (TIC, Stassun et al. 2018, 2019) gives a flux contamination ratio (i.e., ratio of the total contaminant flux within 210″ to HIP 65426’s flux) of only 0.03. Overall, it is likely that significant signals in our TESS data correspond to HIP 65426.
@@ 3.1 High-Frequency Scuti Pulsations in HIP 65426
We calculated the amplitude spectrum of each TESS sector time series (Figure 1abc) using lightkurve (Lightkurve Collaboration et al., 2018) and identified several significant frequencies spanning 3–130 cycles day (Figure 1d). For the high frequencies, we used SigSpec (Reegen, 2007) to iteratively fit sine waves to the data above 10 cycles day, and up to a spectral significance threshold of 10. This resulted in 13 significant pulsation frequencies for Sector 11 and 14 significant pulsation frequencies for Sectors 38 and 64. The pulsation frequencies span 28–131 cycles day, consistent with pressure modes of Scuti stars (e.g., Kurtz, 2022), and they are tabulated in Table 2.
@@
We attempted an asteroseismic mode identification for HIP 65426 by constraining the large frequency seperation (). We used the echelle package (Hey & Ball, 2020) to iterate through a range of trial values and visually inspected the échelle diagrams for regular vertical sequences following Bedding et al. (2020). We arrived at a possible value of 7.230.02 cycles day (Figure 2b). This is consistent within the range that Bedding et al. (2020) found for their sample of high-frequency Scuti stars, but the resulting ridge patterns of HIP 65426 are too complex to confidently identify the majority of the pulsation modes. The rapid stellar rotation period (indicated by and our analysis in Section 3.2) greatly contributes to the complication (e.g., Reese, 2022; Aerts & Tkachenko, 2023, and references therein). Detailed asteroseismic modeling, which may yield an independent age estimate, is beyond the scope of this study.
@@ 3.2 Rapid Differential Rotation in HIP 65426
Inspection of the cycles day region of the amplitude spectrum reveals several significant signals that are well-separated from the pressure modes (Figure 3). We identify two rotation frequencies, =3.06 cycles day and = 3.60 cycles day. is detected in all three TESS sectors whereas is detected is only Sector 64. Gaussian functions were fit to the highest peak of each frequency and the best-fitting standard deviation is the reported uncertainty (Table 1). In the case of , our adopted value is the average of the three separate detections from each TESS sector. This yielded =3.060.03 cycles day (7.850.08 hr) and =3.600.02 cycles day (6.670.04 hr).
We associated these signals with the stellar rotation period for the following reasons. Balona (2017) compared Kepler photometric rotation frequencies for a sample of 30 hot (8300–12000 K) stars with their corresponding values and found that they correlated in a manner physically consistent with rotation frequencies. We overplotted and compared with the value of HIP 65426 (280 km s, Section 3.3) along with the sample presented in Balona (2017) in Figure 4. Both follow the expected trend, which validates our interpretation of the signals as tracers of the stellar rotation frequency. We considered if HIP 65426 could be a hybrid pulsator showing gravity modes in addition to pressure modes (e.g., Grigahcène et al., 2010). However, low-frequency gravity modes often appear in groups of several dense peaks (e.g., Li et al., 2020), as opposed to more narrow signals like in the case of HIP 65426. Moreover, the TESS amplitude spectra show signal at and (Figure 3) consistent with rotation harmonics and an untypical feature for pulsations (e.g., Uytterhoeven et al., 2011).
@@ 3.3 HIP 65426 Stellar Inclination
Chauvin et al. (2017) reported an isochrone-derived stellar radius of , which is consistent with the TIC value of (Stassun et al., 2018, 2019) derived using the Stefan-Boltzmann law. As an input radius for HIP 65426, we used the mean of these two literature values. Similarly for the input , we used the mean of two reported values measured using HARPS spectroscopy: 2999 km s (Chauvin et al., 2017) and 2612 km s (Petrus et al., 2021). Conservatively, we opted to keep the larger of the two reported statistical uncertainties, i.e. 2809 km s. For the input rotation period, we took the mean of our two TESS rotation periods from Section 3.2 and again conservatively opted to keep the larger of the two uncertainties, i.e., 7.260.08 hr.
The resulting posterior for is shown in Figure 5. Here we have assumed a prograde orbit of HIP 65426 b relative to its host star, whereby the planet orbits clockwise on the sky (i.e., is inclined at , Section 3.4), and thus the inclination angle parameter space spans [90,180]. The posterior median and 68% credible interval is , with a maximum a posteriori of 107. This indicates that HIP 65426 is most likely oriented somewhat edge-on, consistent with a priori expectation from the literature measurements. The 95% and 99.7% credible intervals span 91–118 and 90–122, respectively.
@@ 3.4 HIP 65426 b Orbital Inclination
The most important result for the purposes of our study is the orbital inclination posterior (), which is shown in Figure 5. The median and 68% credible interval is ; the 95% credible interval spans 103–139. Our inclination posterior is consistent within the 68% uncertainties to those of recent studies (e.g., Carter et al., 2023; Do Ó et al., 2023; Blunt et al., 2023), despite minor differences in the choices of input astrometry and priors.
@@ 4.1 A Lack of Evidence for Misalignment
The inclination posteriors for HIP 65426 b and its host star are compared in Figure 5. They are consistent within their 68% uncertainties, which we interpret as a lack of significant evidence for misalignment. In the absence of knowledge of the stellar rotation axis orientation, the best available proxy for the true star-planet obliquity is the inclination difference of HIP 65426 b and its host star, , which represents a minimum inclination difference with respect to the true obliquity. We place 68% and 95% upper limits of and , respectively, for the HIP 65426 system.
@@ 5 Conclusion
We analyzed TESS time-series photometry of the directly-imaged exoplanet host star HIP 65426. We detected several pulsation modes consistent with classification as a high-frequency Scuti pulsator. We find a preliminary estimate of 7.230.02 cycles day, consistent with other young Scuti stars (Bedding et al., 2020). We also used the nondetection of pulsation timing variations to estimate a 95% dynamical mass upper limit of 12.8 for HIP 65426 b, independently supporting its planetary mass.
From the TESS data we also detected rapid photometric rotation frequencies of 7.850.08 hr and 6.670.04 hr that we interpreted as evidence of latitudinal differential rotation in HIP 65426. Using published radius and constraints together with our TESS rotation frequencies, we jointly measured the inclination of HIP 65426 to . This near edge-on inclination of the host star is consistent with HIP 65426 b’s near edge-on inclination of that we determined from our orbit fit of the available relative astrometry. We thus found no significant evidence for spin-orbit misalignment in the HIP 65426 system. This finding further supports an emerging trend of preferential alignment between imaged long-period giant planets and their host stars (Figure 7), which is similar to the emerging trend for debris disks and dissimilar to that for imaged brown dwarfs.
@@
Acknowledgements. AGS thanks Brendan P. Bowler and Sarah Blunt for collaborative discussions that helped expedite this work, as well as Mitchell T. Dennis for assistance with operating the HPC clusters Mana and Koa.

### v2

@@ Abstract
HIP 65426 hosts a young giant planet that has become the first exoplanet directly imaged with JWST. Using time-series photometry from the Transiting Exoplanet Survey Satellite (TESS), we classify HIP 65426 as a high-frequency Scuti pulsator with a possible large frequency separation of 7.230.02 cycles day-1. We check the TESS data for pulsation timing variations and use the nondetection to estimate a 95% dynamical mass upper limit of 12.8 for HIP 65426 b. We also identify a low-frequency region of signal that we interpret as stellar latitudinal differential rotation with two rapid periods of 7.850.08 hr and 6.670.04 hr. We use our TESS rotation periods together with published values of radius and to jointly measure the inclination of HIP 65426 to ∘. Our stellar inclination is consistent with the orbital inclination of HIP 65426 b (∘) at the percent level based on our orbit fit using published relative astrometry. The lack of significant evidence for spin-orbit misalignment in the HIP 65426 system supports an emerging trend consistent with preferential alignment between imaged long-period giant planets and their host stars.
@@ 1 Introduction
HIP 65426 (HD 116434, TIC 438702139) is a nearby (107.50.4 pc, Gaia Collaboration et al. 2016, 2021) A2 V star (Houk, 1978) that hosts a directly imaged giant planet (HIP 65426 b, Chauvin et al., 2017). The host star is a member of the Lower Centaurus-Crux (LCC) young moving group (de Zeeuw et al., 1999; Luhman, 2022), with an age of 10–23 Myr (e.g., Mamajek et al., 2002; Sartori et al., 2003; Pecaut et al., 2012; Song et al., 2012). Independent of its membership in LCC, Chauvin et al. (2017) derived an age of 144 Myr for HIP 65426 primarily using isochrone fitting. HIP 65426 has several astrophysically interesting qualities in addition to being an exoplanet host star. Its (261–299 km s-1, Chauvin et al. 2017; Petrus et al. 2021) indicates that it is rapidly rotating. Hints of stellar pulsations have been identified in its radial velocity variations (e.g., Chauvin et al., 2017; Grandjean et al., 2020; Petrus et al., 2021), but HIP 65426’s variability status has not yet been formally confirmed and classified. A detection of high-frequency Scuti pulsations would independently confirm HIP 65426’s youth (Bedding et al., 2020) and might even enable an asteroseismic age to be measured (e.g., Murphy et al., 2021; Steindl et al., 2022).
The giant planet HIP 65426 b orbits at a semi-major axis of 62–120 au and inclined 99–112∘ to the line of sight based on orbit fitting analyses (Cheetham et al., 2019; Bowler et al., 2020; Carter et al., 2023; Do Ó et al., 2023; Blunt et al., 2023). Mass estimates of HIP 65426 b, derived from evolution models and from spectral fitting, span 6–11 (Chauvin et al., 2017; Cheetham et al., 2019; Marleau et al., 2019; Petrus et al., 2021; Carter et al., 2023). HIP 65426 was recently observed with NIRCam and MIRI as part of the JWST (Gardner et al., 2006, 2023) Early Release Science Program (Hinkley et al., 2022), resulting in HIP 65426 b becoming the first direct detection of an exoplanet at wavelengths beyond 5 m (Carter et al., 2023).
@@ 2 TESS Observations
HIP 65426 was observed by TESS in 2-minute cadence mode for Sectors 11, 38, and 64. The gaps between these sectors are 2 yr. Sector 11 observations spanned UTC 23 April 2019–20 May 2019, Sector 38 spanned UTC 29 April 2021–26 May 2021, and Sector 64 spanned UTC 06 April 2023–04 May 2023. We used lightkurve (Lightkurve Collaboration et al., 2018) to download PDC-SAP light curves processed by SPOC (Smith et al., 2012; Stumpe et al., 2012, 2014; Jenkins et al., 2016). We removed outlying photometry from each TESS sector via a sigma-clipping of five standard deviations from the median flux value, which is a standard practice to remove outliers potentially caused by instrumental artifacts that were not captured through standard quality flags. This resulted in only 21 cadences () being removed, resulting in 16,392 cadences for Sector 11, 18,486 cadences for Sector 38, and 18,849 cadences for Sector 64.
We assessed our TESS data for any significant sources of contamination. We began by inspecting a region within a radius of 80″ from HIP 65426, which is an approximate size of the default aperture used for our light curves. We identified no Gaia sources with comparable brightness to HIP 65426, which has an apparent Gaia -band magnitude of 7.00 (Gaia Collaboration et al., 2016, 2023). The five brightest Gaia sources after HIP 65426 in this region (spanning separations of 17–47″) have respective apparent -band magnitudes of 11.98, 13.28, 14.12, 14.59, and 15.05. Their respective colors are 1.72, 0.72, 1.41, 1.47, and 1.32. HIP 65426 has no known stellar companions (Kouwenhoven et al., 2005; Chauvin et al., 2017; Petrus et al., 2021). Furthermore, the TESS Input Catalog (TIC, Stassun et al. 2018, 2019) gives a flux contamination ratio (i.e., ratio of the total contaminant flux within 210″ to HIP 65426’s flux) of only 0.03 (See also Section 3.2). Overall, it is likely that significant signals in our TESS data correspond to HIP 65426.
@@ 3.1 High-Frequency Scuti Pulsations in HIP 65426
We calculated the amplitude spectrum of each TESS sector time series (Figure 1abc) using lightkurve (Lightkurve Collaboration et al., 2018) and identified several significant frequencies spanning 3–130 cycles day-1 (Figure 1d). For the high frequencies, we used SigSpec (Reegen, 2007) to iteratively fit sine waves to the data above 10 cycles day-1, and up to a spectral significance threshold of 10 (following, e.g., Gruberbauer et al., 2007; Zwintz et al., 2009; Sepulveda et al., 2022). This resulted in 13 significant pulsation frequencies for Sector 11 and 14 significant pulsation frequencies for Sectors 38 and 64. The pulsation frequencies span 28–131 cycles day-1, consistent with pressure modes of Scuti stars (e.g., Kurtz, 2022), and they are tabulated in Table 2.
@@
We attempted an asteroseismic mode identification for HIP 65426 by constraining the large frequency separation (), which is a parameter that can also constrain the mean stellar density (Aerts et al., 2010). We used the echelle package (Hey & Ball, 2020) to iterate through a range of trial values (5–9 cycles day-1, which generously encompasses typical values, e.g., Bedding et al. 2020; Murphy et al. 2023) and visually inspected the échelle diagrams for regular vertical sequences following Bedding et al. (2020). We arrived at a possible value of 7.230.02 cycles day-1 (Figure 2b). This is consistent within the range that Bedding et al. (2020) found for their sample of high-frequency Scuti stars, but the resulting ridge patterns of HIP 65426 are too complex to confidently identify the majority of the pulsation modes. The rapid stellar rotation period (indicated by and our analysis in Section 3.2) greatly contributes to the complication (e.g., Reese, 2022; Aerts & Tkachenko, 2023, and references therein). Detailed asteroseismic modeling, which may yield an independent age estimate, is beyond the scope of this study.
@@ 3.2 Rapid Differential Rotation in HIP 65426
Inspection of the cycles day-1 region of the amplitude spectrum reveals several significant signals that are well-separated from the pressure modes (Figure 3). We identify two rotation frequencies, =3.06 cycles day-1 and = 3.60 cycles day-1. is detected in all three TESS sectors whereas is detected is only Sector 64. Gaussian functions were fit to the highest peak of each frequency and the best-fitting standard deviation is the reported uncertainty (Table 1). In the case of , our adopted value is the average of the three separate detections from each TESS sector. This yielded =3.060.03 cycles day-1 (7.850.08 hr) and =3.600.02 cycles day-1 (6.670.04 hr). We note that is similar to our estimated value for . This is likely a coincidence, but supports the conclusion that rapid rotation will complicate the mode identification for this star.
We associated these signals with the stellar rotation period for the following reasons. Balona (2017) compared Kepler photometric rotation frequencies for a sample of 30 hot (8300–12000 K) stars with their corresponding values and found that they correlated in a manner physically consistent with rotation frequencies. We overplotted and compared with the value of HIP 65426 (280 km s-1, Section 3.3) along with the sample presented in Balona (2017) in Figure 4. Both follow the expected trend, which validates our interpretation of the signals as tracers of the stellar rotation frequency. We considered if HIP 65426 could be a hybrid pulsator showing gravity modes in addition to pressure modes (e.g., Grigahcène et al., 2010). However, low-frequency gravity modes often appear in groups of several dense peaks (e.g., Li et al., 2020), as opposed to more narrow signals like in the case of HIP 65426. Moreover, the TESS amplitude spectra show signal at and (Figure 3) consistent with rotation harmonics and an untypical feature for pulsations (e.g., Uytterhoeven et al., 2011). The peak at 6.6 cycles day-1 is associated with rotation, but we do not include it in Table 1 due to the absence of a subharmonic. This is not uncommon for rotational modulation (e.g. Kawaler, 2021) and could be caused by multiple spots located at different longitudes. The same is likely true for the frequency in Sector 38 without .
@@ 3.3 HIP 65426 Stellar Inclination
Chauvin et al. (2017) reported an isochrone-derived stellar radius of , which is consistent with the TIC value of (Stassun et al., 2018, 2019) derived using the Stefan-Boltzmann law. As an input radius for HIP 65426, we used the mean of these two literature values. We note however that rapidly rotating intermediate mass stars are better described as oblate spheroids than spheres (e.g., van Belle et al., 2001; Domiciano de Souza et al., 2003; McAlister et al., 2005; Monnier et al., 2007). Therefore, HIP 65426 is likely not well described by a single radius but rather by a radius gradient that increases from polar to equatorial latitudes. HIP 65426 is too distant (i.e., its angular size is too small) to resolve with optical/infrared interferometry, a technique that can directly constrain the polar () and equatorial () radii for nearby stars with large projected angular sizes (e.g., van Belle, 2012). To account for systematic uncertainty in our input radius due to unresolved oblateness, we consider and of Altair, a nearby Scuti star whose surface has been interferometrically resolved and who is of similar mass and rotational velocity as HIP 65426 (e.g., Buzasi et al., 2005; Monnier et al., 2007; Bouchaud et al., 2020). Using = 1.661 and = 2.022 (Monnier et al., 2007) we calculate a fractional uncerainty of , which we adopt for our input value for HIP 65426 (i.e., .)
For the input , we used the mean of two reported values measured using HARPS spectroscopy: 2999 km s-1 (Chauvin et al., 2017) and 2612 km s-1 (Petrus et al., 2021). For the uncertainty, we used the standard error on the mean of the HARPS values, yielding 28013 km s-1. We note that in the presence of latitudinal differential rotation, the spectroscopic (derived without accounting for differential rotation) will represent a weighted average of the set of latitudinal velocities and not soley represent the equatorial velocity. This is supported by interferometric imaging of the nearby A-star Altair, which has a of 241 km s-1 from detailed modeling (Monnier et al., 2007) and a from methods that ignore differential rotation of 216 km s-1 (mean Altair from: Bernacca & Perinotto, 1970; Abt & Morrell, 1995; Royer et al., 2002; Schröder et al., 2009). Detailed modeling of the differential rotation (and its effect on line broadening) in HIP 65426 is beyond the scope of this paper.
For the input rotation period, we took the mean of our two TESS rotation periods from Section 3.2. By taking the mean we appropriately match the weighted average velocity implied by the HARPS measurements. We opted to keep the larger of the two individual period uncertainties for our input mean period , i.e., 7.260.08 hr.
The resulting posterior for is shown in Figure 5. Here we have assumed a prograde orbit of HIP 65426 b relative to its host star, whereby the planet orbits clockwise on the sky (i.e., is inclined at , Section 3.4), and thus the inclination angle parameter space spans [90,180]∘. The posterior median and 68% credible interval is ∘, with a maximum a posteriori of 107∘. This indicates that HIP 65426 is most likely oriented somewhat edge-on, consistent with a priori expectation from the literature measurements. The 95% and 99.7% credible intervals span 91–127∘ and 90–133∘, respectively.
@@ 3.4 HIP 65426 b Orbital Inclination
The most important result for the purposes of our study is the orbital inclination posterior (), which is shown in Figure 5. The median and 68% credible interval is ∘; the 95% credible interval spans 103–139∘. Our inclination posterior is consistent within the 68% uncertainties to those of recent studies (e.g., Carter et al., 2023; Do Ó et al., 2023; Blunt et al., 2023), despite minor differences in the choices of input astrometry and priors.
@@ 4.1 A Lack of Evidence for Misalignment
The inclination posteriors for HIP 65426 b and its host star are compared in Figure 5. They are consistent within their 68% uncertainties, which we interpret as a lack of significant evidence for misalignment. In the absence of knowledge of the stellar rotation axis orientation, the best available proxy for the true star-planet obliquity is the inclination difference of HIP 65426 b and its host star, , which represents a minimum inclination difference with respect to the true obliquity. We place 68% and 95% upper limits of ∘ and ∘, respectively, for the HIP 65426 system.
@@ 5 Conclusion
We analyzed TESS time-series photometry of the directly-imaged exoplanet host star HIP 65426. We detected several pulsation modes consistent with classification as a high-frequency Scuti pulsator. We find a preliminary estimate of 7.230.02 cycles day-1, consistent with other young Scuti stars (Bedding et al., 2020). We also used the nondetection of pulsation timing variations to estimate a 95% dynamical mass upper limit of 12.8 for HIP 65426 b, independently supporting its planetary mass.
From the TESS data we also detected rapid photometric rotation frequencies of 7.850.08 hr and 6.670.04 hr that we interpreted as evidence of latitudinal differential rotation in HIP 65426. Using published radius and constraints together with our TESS rotation frequencies, we jointly measured the inclination of HIP 65426 to ∘. This near edge-on inclination of the host star is consistent with HIP 65426 b’s near edge-on inclination of ∘ that we determined from our orbit fit of the available relative astrometry. We thus found no significant evidence for spin-orbit misalignment in the HIP 65426 system. This finding further supports an emerging trend of preferential alignment between imaged long-period giant planets and their host stars (Figure 7), which is similar to the emerging trend for debris disks and dissimilar to that for imaged brown dwarfs.
@@
Acknowledgments. AGS thanks Brendan P. Bowler and Sarah Blunt for collaborative discussions that helped expedite this work, as well as Mitchell T. Dennis for assistance with operating the HPC clusters Mana and Koa.